AnimeGhoul8863:

An expression is shown below: 3pf2 − 21p2f + 6pf − 42p2 Part A: Rewrite the expression so the GCF is factored completely. (4 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (6 points)

AnimeGhoul8863:

\[3pf^2-21p^2f+6pf-42p^2\]

Spectre:

Are you solving it, or am I?

AnimeGhoul8863:

i need help solving it i was just making the equation easier to read

Spectre:

So, do you have any idea how to do this?

Pixel:

@Shadow

AnimeGhoul8863:

3p (f-7p) (f+2) and the GCF is 3p

Shadow:

You can check your work by throwing that in wolfram and seeing if it comes out to the expression you started with. I’m on mobile so I can’t be of much help, sorry.

AnimeGhoul8863:

@Spectre

Spectre:

I got (3pf - 21p²)(f + 2)

AnimeGhoul8863:

hmmmmmmm

AnimeGhoul8863:

@Spectre can u help me im confused

Spectre:

3pf² - 21p²f + 6pf - 42p² = 3pf² + 6pf - 21p²f - 42p² = 3pf(f +2) - 21p²(f + 2) = (3pf - 21p²)(f +0 2) Answer: (3pf - 21p²)(f + 2)

AnimeGhoul8863:

so if ^^^^ is part a what would part B be

AnimeGhoul8863:

so do this Part A: (3pf - 21p²)(f + 2) Part B: 3pf² - 21p²f + 6pf - 42p² = 3pf² + 6pf - 21p²f - 42p² = 3pf(f +2) - 21p²(f + 2) = (3pf - 21p²)(f +0 2) Factored GCF: (3pf - 21p²)(f + 2)

Shadow:

Looks right to me

Spectre:

Yes.